# Transfer of quantum game strategies

### v1

@@ 3. Strongly stochastic operator matrices
[…]
We also will consider the space
$\mathcal{S}_{X,A}:={\rm span}\{e_{x,x,a,a}:\;x\in X,a\in A\},$
viewed as an operator subsystem inside $\mathcal{T}_{X,A}$.
[…]
@@
Setting $\tilde{P}=(P_{xx^{\prime},yy^{\prime}}^{aa^{\prime},bb^{\prime}})_{xx^{\prime},yy^{\prime},aa^{\prime},bb^{\prime}}\in M_{XYAB}\otimes\mathcal{B}(H)$, one may easily check that $\tilde{P}$ is strongly stochastic over $(XY,AB)$ with ${\rm Tr}_{B}(L_{\sigma_{Y}}(\tilde{P}))$ and ${\rm Tr}_{A}(L_{\sigma_{X}}(\tilde{P}))$ classical stochastic operator matrices (as introduced in [38, Section 3]) for each (pure) state $\sigma_{X}\in M_{X}$ and $\sigma_{Y}\in M_{Y}$. Thus, we may think of strongly stochastic operator matrices over $(XY,AB)$ as generalizations of NS operator matrices, whose “marginal” stochastic operators are no longer necessarily classical.
[…]
(9) $P_{xx^{\prime},yy^{\prime}}^{aa^{\prime},bb^{\prime}}=V^{*}E_{xx^{\prime},aa^{\prime}}F_{yy^{\prime},bb^{\prime}}V,\;\;\;\;x,x^{\prime}\in X,y,y^{\prime}\in Y,a,a^{\prime}\in A,b,b^{\prime}\in B.$
In this section and the next, we wish to link strongly stochastic operator matrices to particular subclasses of QNS correlations, and the (multi-variate) tensor product of operator systems. In order to do so, we first briefly recall some facts about operator systems and related constructions. If $\mathcal{S}$ and $\mathcal{T}$ are operator systems, we call $\mathcal{S}$ and $\mathcal{T}$ completely order isomorphic and write $\mathcal{S}\cong_{\rm c.o.i}\mathcal{T}$ if there exists a unital completely positive bijection $\phi:\mathcal{S}\rightarrow\mathcal{T}$ with completely positive inverse. We write $\mathcal{S}\subseteq_{\rm c.o.i}\mathcal{T}$ if $\mathcal{S}\subseteq\mathcal{T}$ and the inclusion map $\iota:\mathcal{S}\hookrightarrow\mathcal{T}$ is a complete order isomorphism onto its range. The three main types of operator system tensor products that will be used in the sequel are given as follows:
  • (i)
  • (ii)
  • (iii)
More details about each tensor product may be found in [26]; the construction of the multivariate tensor product of each type, with explicit descriptions of their matricial cones, may be found in [19, Section 7].
We next recall the notion of a coproduct of operator systems: if $\mathcal{S}$ and $\mathcal{T}$ are two operator systems, their coproduct $\mathcal{S}\oplus_{1}\mathcal{T}$ is the unique (up to isomorphism) operator system equipped with complete order embeddings $\iota_{\mathcal{S}}:\mathcal{S}\rightarrow\mathcal{S}\oplus_{1}\mathcal{T}$ and $\iota_{\mathcal{T}}:\mathcal{T}\rightarrow\mathcal{S}\oplus_{1}\mathcal{T}$ which satisfies the following unversal property: For each ucp map $\phi:\mathcal{S}\rightarrow\mathcal{R}$ and $\psi:\mathcal{T}\rightarrow\mathcal{R}$, where $\mathcal{R}$ is an operator system, there exists a unique ucp map $\varphi:\mathcal{S}\oplus_{1}\mathcal{T}\rightarrow\mathcal{R}$ such that $\varphi(\iota_{\mathcal{S}}(s))=\phi(s)$ and $\varphi(\iota_{\mathcal{T}}(t))=\psi(t)$ for every $s\in\mathcal{S},t\in\mathcal{T}$. For more properties of the coproduct of operator systems, we refer the reader to [24, Section 8].
For the remainder of this section, to help distinguish between operator systems $\mathcal{T}_{X,A}$ and $\mathcal{T}_{Y,B}$, we will denote the canonical generators of $\mathcal{T}_{X,A}$ by $e_{x,x^{\prime},a,a^{\prime}},x,x^{\prime}\in X,a,a^{\prime}\in A$ and $\mathcal{T}_{Y,B}$ by $f_{y,y^{\prime},b,b^{\prime}},y,y^{\prime}\in Y,b,b^{\prime}\in B$.
Proposition 3.3.
Proof.
Let $K$ be a Hilbert space, $V:H\rightarrow K$ be an isometry, and
be mutually commuting stochastic operator matrices on $K$ satisfying (9). Define linear maps $\phi:\mathcal{T}_{X,A}\rightarrow\mathcal{B}(K)$ (resp. $\psi:\mathcal{T}_{Y,B}\rightarrow\mathcal{B}(K)$) by $\phi(e_{x,x^{\prime},a,a^{\prime}})=E_{xx^{\prime},aa^{\prime}}$ (resp. $\psi(f_{y,y^{\prime},b,b^{\prime}})=F_{yy^{\prime},bb^{\prime}}$). By [38, Theorem 5.2], both $\phi$ and $\psi$ are unital and completely positive. By the definition of the commuting tensor product of operator spaces, the map $\phi\cdot\psi:\mathcal{T}_{X,A}\otimes_{\rm c}\mathcal{T}_{Y,B}\rightarrow\mathcal{B}(K)$ given by $(\phi\cdot\psi)(u\otimes v)=\phi(u)\psi(v)$ is (unital and) completely positive as well. Set
Conversely, suppose that $\gamma:\mathcal{T}_{X,A}\otimes_{\rm c}\mathcal{T}_{Y,B}\rightarrow\mathcal{B}(H)$ is a unital completely positive map. By [38, Corollary 5.3] and [26, Theorem 6.4], $\mathcal{T}_{X,A}\otimes_{\rm c}\mathcal{T}_{Y,B}\subseteq_{\rm c.o.i}\mathcal{C}_{X,A}\otimes_{\rm max}\mathcal{C}_{Y,B}$. Apply Arveson’s Extension Theorem to obtain a completely positive map $\tilde{\gamma}:\mathcal{C}_{X,A}\otimes_{\rm max}\mathcal{C}_{Y,B}\rightarrow\mathcal{B}(H)$ extending $\gamma$. If we then apply Stinespring’s Theorem, we may write
Remark 3.4.
By Remark 3.2 and [19, Remark 5.3], there exist strongly stochastic operator matrices which are not dilatable.
[…]
(10) $P_{xx^{\prime},yy^{\prime}}^{aa^{\prime},bb^{\prime}}=V^{*}(E_{xx^{\prime},aa^{\prime}}\otimes F_{yy^{\prime},bb^{\prime}})V,$
@@
(11) $\Gamma(aa^{\prime},bb^{\prime}|xx^{\prime},yy^{\prime})=\langle P_{xx^{\prime},yy^{\prime}}^{aa^{\prime},bb^{\prime}}\xi,\xi\rangle,$
@@
First, assume that $\Gamma\in\mathcal{Q}_{\rm loc}$ is a convex combination $\Gamma=\sum_{j=1}^{k}\lambda_{j}\Phi^{(j)}\otimes\Psi^{(j)}$, where $\Phi^{(j)}:M_{X}\rightarrow M_{A}$ and $\Psi^{(j)}:M_{Y}\rightarrow M_{B}$ are quantum channels, $j=1,\ldots,k$. By the comment before [38, Remark 3.2], $\Phi^{(j)}$ (resp. $\Psi^{(j)}$) is of the form $\Gamma_{E^{(j)},1}$ (resp. $\Gamma_{F^{(j)},1}$) for some stochastic operator matrix $E^{(j)}\in M_{X}\otimes M_{A}\otimes\mathcal{B}(\mathbb{C})$ (resp. $F^{(j)}\in M_{Y}\otimes M_{B}\otimes\mathcal{B}(\mathbb{C})$) for $j=1,\ldots,k$. For each $x,x^{\prime}\in X,a,a^{\prime}\in A$ define the matrix $E_{x,x^{\prime}}^{a,a^{\prime}}=(E_{x,x^{\prime},a,a^{\prime}}^{(j)})_{j=1}^{k}\in\mathcal{D}_{k}$. Define the corresponding matrices $F_{y,y^{\prime}}^{b,b^{\prime}}$ in $\mathcal{D}_{k}$ for each $y,y^{\prime}\in Y,b,b^{\prime}\in B$. We note that the family
@@
Let $\mathcal{A}_{X}^{A}$ (resp. $\mathcal{A}_{Y}^{B}$) be the abelian ${\rm C}^{*}$-algebra generated by $\{E_{x,x^{\prime},a,a^{\prime}}:\;x,x^{\prime}\in X,a,a^{\prime}\in A\}$ (resp. $\{F_{y,y^{\prime},b,b^{\prime}}:\;y,y^{\prime}\in Y,b,b^{\prime}\in B\}$). Let $s:\mathcal{A}_{X}^{A}\otimes_{\rm max}\mathcal{A}_{Y}^{B}\rightarrow\mathbb{C}$ be the state given by $s(S\otimes T)=\langle ST\xi,\xi\rangle$. By nuclearity of abelian ${\rm C}^{*}$-algebras, we may consider $s$ as a state on $\mathcal{A}_{X}^{A}\otimes_{\rm min}\mathcal{A}_{Y}^{B}$. Using the identification $\mathcal{A}_{X}^{A}=C(\Omega_{1}),\mathcal{A}_{Y}^{B}=C(\Omega_{2})$ for compact Hausdorff spaces $\Omega_{1},\Omega_{2}$, we view $s$ as a Borel probability measure $\mu$ on the product topological space $\Omega_{1}\times\Omega_{2}$. For each $\omega_{1}\in\Omega_{1}$ (resp. $\omega_{2}\in\Omega_{2}$), let $\Phi(\omega_{1}):M_{X}\rightarrow M_{A}$ (resp. $\Psi(\omega_{2}):M_{Y}\rightarrow M_{B}$) be the quantum channel given by $\Phi(\omega_{1})(\epsilon_{x,x^{\prime}})=(E_{x,x^{\prime},a,a^{\prime}}(\omega_{1}))_{a,a^{\prime}}$ (resp. $\Psi(\omega_{2})(\epsilon_{y,y^{\prime}})=(F_{y,y^{\prime},b,b^{\prime}}(\omega_{2}))_{b,b^{\prime}}$). We then have
@@
[…]
@@
If we approximate measure $\mu$ using convex combinations of product measures $\mu_{1}\times\mu_{2}$ (where $\mu_{1}\in{\rm M}(\Omega_{1}),\mu_{2}\in{\rm M}(\Omega_{2})$) we see we can approximate $\Gamma$ using channels which may be written as convex combinations $\tilde{\Gamma}=\sum_{j=1}^{k}\lambda_{j}\Phi_{j}\otimes\Psi_{j}$, where $\Phi_{j}:M_{X}\rightarrow M_{A}$ and $\Psi_{j}:M_{Y}\rightarrow M_{B}$ are quantum channels, $j=1,\ldots,k$. By the Carathéodory Theorem and compactness, use a similar argument as in [38, Remark 4.10] to conclude that $\Gamma$ itself is of that form. Thus, $\Gamma\in\mathcal{Q}_{\rm loc}$. ∎
3.1. Strongly bistochastic operator matrices
For the remainder of this section, assume that $X=A$ and $Y=B$. For notational ease, we will continue to refer to $X$ and $A$ (resp. $Y$ and $B$) as distinct entities, even though they are copies of the same set.
A positive operator $P=(P_{xx^{\prime},yy^{\prime}}^{aa^{\prime},bb^{\prime}})_{xx^{\prime},yy^{\prime},aa^{\prime},bb^{\prime}}\in M_{XYAB}\otimes\mathcal{B}(H)$ will be called a strongly bistochastic operator matrix if it is a strongly stochastic operator matrix over $(XY,AB)$, with ${\rm Tr}_{B}(L_{\sigma_{Y}}(P))$ and ${\rm Tr}_{A}(L_{\sigma_{X}}(P))$ bistochastic operator matrices for each state $\sigma_{X}\in M_{X},\sigma_{Y}\in M_{Y}$.
@@ 3.1. Strongly bistochastic operator matrices
A strongly bistochastic operator matrix $P=(P_{xx^{\prime},yy^{\prime}}^{aa^{\prime},bb^{\prime}})_{xx^{\prime},yy^{\prime},aa^{\prime},bb^{\prime}}$ is called dilatable if there exists a Hilbert space $K$, an isometry $V:H\rightarrow K$, and bistochastic operator matrices $(E_{xx^{\prime},aa^{\prime}})_{x,x^{\prime}\in X,a,a^{\prime}\in A}$ (resp. $(F_{yy^{\prime},bb^{\prime}})_{y,y^{\prime}\in Y,b,b^{\prime}\in B}$) in $M_{XA}\otimes\mathcal{B}(K)$ (resp. $M_{YB}\otimes\mathcal{B}(K)$) such that $E_{xx^{\prime},aa^{\prime}}F_{yy^{\prime},bb^{\prime}}=F_{yy^{\prime},bb^{\prime}}E_{xx^{\prime},aa^{\prime}}$ and which satisfy relations (9) for all $x,x^{\prime}\in X,a,a^{\prime}\in A,y,y^{\prime}\in Y,b,b^{\prime}\in B$. Locally dilatable and quantum dilatable strongly bistochastic operator matrices are defined exactly as with strongly stochastic operator matrices, but using bistochastic operator matrices in place of stochastic operator matrices.
To close this section, we recall the universal operator system corresponding to bistochastic operator matrices (first introduced in [7, Section 3]) which will be needed to show an analogous result to Proposition 3.3.
Let $\mathcal{V}_{X}$ be the universal TRO generated by the entries $v_{a,x}$ of a block operator bi-isometry $V=(v_{a,x})_{a\in A,x\in X}$ (see [7]). Let $\mathcal{C}_{X}$ be the right ${\rm C}^{*}$-algebra of $\mathcal{V}_{X}$, set $e_{x,x^{\prime},a,a^{\prime}}=v_{a,x}^{*}v_{a,x}$ (where the latter is considered as an element of $\mathcal{C}_{X}$) and
be viewed as an operator system in $\mathcal{C}_{X}$. Furthermore, we let
$\mathcal{S}_{X}:={\rm span}\{e_{x,x,a,a}:\;x\in X,a\in A\}$
be viewed as an operator subsystem of $\mathcal{T}_{X}$.
Proposition 3.6.
Proof.
Using [7, Theorem 3.4] and [26, Theorem 6.4], argue exactly as in Proposition 3.3 but with bistochastic operator matrices in place of stochastic operator matrices and $\mathcal{T}_{X}$ (resp. $\mathcal{T}_{Y}$) in place of $\mathcal{T}_{X,A}$ (resp. $\mathcal{T}_{Y,B}$). ∎
4. SQNS correlations

### v2

@@ 3. Strongly stochastic operator matrices
Let $X,Y,A$ and $B$ be finite sets, and $H$ be a Hilbert space. In the sequel, to simplify notation we will abbreviate an ordered pair $(x,y)\in X\times Y$ to $xy$. A stochastic operator matrix $P=(P_{xx^{\prime},yy^{\prime}}^{aa^{\prime},bb^{\prime}})_{xx^{\prime},yy^{\prime},aa^{\prime},bb^{\prime}}\in M_{XYAB}\otimes\mathcal{B}(H)$ over $(XY,AB)$ will be called a $\it strongly\;stochastic\;operator\;matrix$} over $(X,Y,A,B)$ if ${\rm Tr}_{B}(L_{\sigma_{Y}}(P))$ (resp. ${\rm Tr}_{A}(L_{\sigma_{X}}(P))$) is a well-defined stochastic operator matrix over $(X,A)$ (resp. $(Y,B)$) and ${\rm Tr}_{B}(L_{\sigma_{Y}}(P))={\rm Tr}_{B}(L_{\sigma_{Y}^{\prime}}(P))$ (resp. ${\rm Tr}_{A}(L_{\sigma_{X}}(P))={\rm Tr}_{A}(L_{\sigma_{X}^{\prime}}(P))$) for each pure state $\sigma_{X},\sigma_{X}^{\prime}\in M_{X}$ and $\sigma_{Y},\sigma_{Y}^{\prime}\in M_{Y}$.
@@
Setting $\tilde{P}=(P_{xx^{\prime},yy^{\prime}}^{aa^{\prime},bb^{\prime}})_{xx^{\prime},yy^{\prime},aa^{\prime},bb^{\prime}}\in M_{XYAB}\otimes\mathcal{B}(H)$, one may easily check that $\tilde{P}$ is strongly stochastic over $(XY,AB)$ with $){\rm Tr}_{B}(L_{\sigma_{Y}}(\tilde{P}$)) and $){\rm Tr}_{A}(L_{\sigma_{X}}(\tilde{P}$)) classical stochastic operator matrices (as introduced in [39, Section 3]) for each (pure) state $\sigma_{X}\in M_{X}$ and $\sigma_{Y}\in M_{Y}$. Thus, we may think of strongly stochastic operator matrices over $(XY,AB)$ as generalizations of NS operator matrices, whose “marginal” stochastic operators are no longer necessarily classical.
A strongly stochastic operator matrix $P=(P_{xx^{\prime},yy^{\prime}}^{aa^{\prime},bb^{\prime}})_{xx^{\prime},yy^{\prime},aa^{\prime},bb^{\prime}}$ is called $\it dilatable$} if there exists a Hilbert space $K$, an isometry $V:H\rightarrow K$ and stochastic operator matrices $(E_{xx^{\prime},aa^{\prime}})_{x,x^{\prime}\in X,a,a^{\prime}\in A}$ (resp. $(F_{yy^{\prime},bb^{\prime}})_{y,y^{\prime}\in Y,b,b^{\prime}\in B}$) in $M_{XA}\otimes\mathcal{B}(K)$ (resp. $M_{YB}\otimes\mathcal{B}(K)$) such that $E_{xx^{\prime},aa^{\prime}}F_{yy^{\prime},bb^{\prime}}=F_{yy^{\prime},bb^{\prime}}E_{xx^{\prime},aa^{\prime}}$ and
$P_{xx^{\prime},yy^{\prime}}^{aa^{\prime},bb^{\prime}}=V^{*}E_{xx^{\prime},aa^{\prime}}F_{yy^{\prime},bb^{\prime}}V,\;\;\;\;x,x^{\prime}\in X,y,y^{\prime}\in Y,a,a^{\prime}\in A,b,b^{\prime}\in B.$ (9)
We will call a strongly stochastic operator matrix $P=(P_{xx^{\prime},yy^{\prime}}^{aa^{\prime},bb^{\prime}})_{xx^{\prime},yy^{\prime},aa^{\prime},bb^{\prime}}$ acting on a Hilbert space $H$ $\it locally\;dilatable$} if there exists a dilation of the form (9), with the additional stipulation that the family $\{E_{xx^{\prime},aa^{\prime}},F_{yy^{\prime},bb^{\prime}}:\;x,x^{\prime}\in X,y,y^{\prime}\in Y,a,a^{\prime}\in A,b,b^{\prime}\in B\}$ is commutative. We will call a strongly stochastic operator matrix $P=(P_{xx^{\prime},yy^{\prime}}^{aa^{\prime},bb^{\prime}})_{xx^{\prime},yy^{\prime},aa^{\prime},bb^{\prime}}$ acting on a Hilbert space $H$ $\it quantum\;dilatable$} if there exist stochastic operator matrices $(E_{xx^{\prime},aa^{\prime}})_{x,x^{\prime}\in X,a,a^{\prime}\in A}$ and $(F_{yy^{\prime},bb^{\prime}})_{y,y^{\prime}\in Y,b,b^{\prime}\in B}$ acting on finite dimensional Hilbert spaces $H_{A}$ and $H_{B}$ respectively, and an isometry $V:H\rightarrow H_{A}\otimes H_{B}$, such that
$P_{xx^{\prime},yy^{\prime}}^{aa^{\prime},bb^{\prime}}=V^{*}(E_{xx^{\prime},aa^{\prime}}\otimes F_{yy^{\prime},bb^{\prime}})V,$ (10)
@@
$\Gamma(aa^{\prime},bb^{\prime}|xx^{\prime},yy^{\prime})=\langle P_{xx^{\prime},yy^{\prime}}^{aa^{\prime},bb^{\prime}}\xi,\xi\rangle,$ (11)
@@
First, assume that $\Gamma\in\mathcal{Q}_{\rm loc}$ is a convex combination $\Gamma=\sum_{j=1}^{k}\lambda_{j}\Phi^{(j)}\otimes\Psi^{(j)}$, where $\Phi^{(j)}:M_{X}\rightarrow M_{A}$ and $\Psi^{(j)}:M_{Y}\rightarrow M_{B}$ are quantum channels, $j=1,\ldots,k$. By the comment before [39, Remark 3.2], $\Phi^{(j)}$ (resp. $\Psi^{(j)}$) is of the form $\Gamma_{E^{(j)},1}$ (resp. $\Gamma_{F^{(j)},1}$) for some stochastic operator matrix $E^{(j)}\in M_{X}\otimes M_{A}\otimes\mathcal{B}(\mathbb{C})$ (resp. $F^{(j)}\in M_{Y}\otimes M_{B}\otimes\mathcal{B}(\mathbb{C})$) for $j=1,\ldots,k$. For each $x,x^{\prime}\in X,a,a^{\prime}\in A$ define the matrix $E_{x,x^{\prime}}^{a,a^{\prime}}=(E_{x,x^{\prime},a,a^{\prime}}^{(j)})_{j=1}^{k}\in\mathcal{D}_{k}$. Define the corresponding matrices $F_{y,y^{\prime}}^{b,b^{\prime}}$ in $\mathcal{D}_{k}$ for each $y,y^{\prime}\in Y,b,b^{\prime}\in B$. We note that the family
@@
[…]
@@
$=$ $\int\limits_{\Omega_{1}\times\Omega_{2}}E_{x,x^{\prime},a,a^{\prime}}(\omega_{1})F_{y,y^{\prime},b,b^{\prime}}(\omega_{2})d\mu(\omega_{1},\omega_{2})$
$=$ $\bigg\langle\int\limits_{\Omega_{1}\times\Omega_{2}}\Phi(\omega_{1})(\epsilon_{x,x^{\prime}})\otimes\Psi(\omega_{2})(\epsilon_{y,y^{\prime}})d\mu(\omega_{1},\omega_{2}),\epsilon_{a,a^{\prime}}\otimes\epsilon_{b,b^{\prime}}\bigg\rangle.$
@@
If we approximate measure $\mu$ using convex combinations of product measures $\mu_{1}\times\mu_{2}$ (where $\mu_{1}\in{\rm M}(\Omega_{1}),\mu_{2}\in{\rm M}(\Omega_{2})$) we see we can approximate $\Gamma$ using channels which may be written as convex combinations $\tilde{\Gamma}=\sum_{j=1}^{k}\lambda_{j}\Phi_{j}\otimes\Psi_{j}$, where $\Phi_{j}:M_{X}\rightarrow M_{A}$ and $\Psi_{j}:M_{Y}\rightarrow M_{B}$ are quantum channels, $j=1,\ldots,k$. By the Carathéodory Theorem and compactness, use a similar argument as in [39, Remark 4.10] to conclude that $\Gamma$ itself is of that form. Thus, $\Gamma\in\mathcal{Q}_{\rm loc}$. ∎
For certain classes of quantum games (for instance, those concerned with the behavior of quantum symmetries of quantum objects as in [7, 13]) we will draw our questions and answer states from the same space; thus, a particular notion of strongly stochastic operator matrix will be required in this context. We assume that $X=A$ and $Y=B$. For notational ease, we will continue to refer to $X$ and $A$ (resp. $Y$ and $B$) as distinct entities, even though they are copies of the same set.
@@
A positive operator $P=(P_{xx^{\prime},yy^{\prime}}^{aa^{\prime},bb^{\prime}})_{xx^{\prime},yy^{\prime},aa^{\prime},bb^{\prime}}\in M_{XYAB}\otimes\mathcal{B}(H)$ will be called a strongly bistochastic operator matrix if it is a strongly stochastic operator matrix over $(XY,AB)$, with ${\rm Tr}_{B}(L_{\sigma_{Y}}(P))$ and ${\rm Tr}_{A}(L_{\sigma_{X}}(P))$ bistochastic operator matrices for each state $\sigma_{X}\in M_{X},\sigma_{Y}\in M_{Y}$. A strongly bistochastic operator matrix $P=(P_{xx^{\prime},yy^{\prime}}^{aa^{\prime},bb^{\prime}})_{xx^{\prime},yy^{\prime},aa^{\prime},bb^{\prime}}$ is called dilatable if there exists a Hilbert space $K$, an isometry $V:H\rightarrow K$, and bistochastic operator matrices $(E_{xx^{\prime},aa^{\prime}})_{x,x^{\prime}\in X,a,a^{\prime}\in A}$ (resp. $(F_{yy^{\prime},bb^{\prime}})_{y,y^{\prime}\in Y,b,b^{\prime}\in B}$) in $M_{XA}\otimes\mathcal{B}(K)$ (resp. $M_{YB}\otimes\mathcal{B}(K)$) such that $E_{xx^{\prime},aa^{\prime}}F_{yy^{\prime},bb^{\prime}}=F_{yy^{\prime},bb^{\prime}}E_{xx^{\prime},aa^{\prime}}$ and which satisfy relations (9) for all $x,x^{\prime}\in X,a,a^{\prime}\in A,y,y^{\prime}\in Y,b,b^{\prime}\in B$. Locally dilatable and quantum dilatable strongly bistochastic operator matrices are defined exactly as with strongly stochastic operator matrices, but using bistochastic operator matrices in place of stochastic operator matrices.
3.1. SQNS correlations
One goal for this work is to develop a method for transferring perfect strategies of one quantum input-output game to another. The framework we use for this purpose is by embedding both quantum games into a single “game-of-games”, whose winning QNS strategies encode the winning information for both simultaneously. The winning QNS strategies are used as simulators in the simulation paradigm, allowing us to take a perfect strategy for the first game and— in conjunction with a winning strategy for the “game-of-games”— construct the desired strategy for the second. In order to adhere to the no-signalling condition for both games, we have to modify the classes of QNS correlations we use as simulators; this leads us to the introduction of a particular sub-class of QNS correlations, which is the focus of this subsection.